# WAVE: Warp-Based View Guidance for Consistent Novel View Synthesis Using a Single Image

### v1

@@ Abstract
Generating high-quality novel views of a scene from a single image requires maintaining structural coherence across different views, referred to as view consistency. While diffusion models have driven advancements in novel view synthesis, they still struggle to preserve spatial continuity across views. Diffusion models have been combined with 3D models to address the issue, but such approaches lack efficiency due to their complex multi-step pipelines. This paper proposes a novel view-consistent image generation method which utilizes diffusion models without additional modules. Our key idea is to enhance diffusion models with a training-free method that enables adaptive attention manipulation and noise reinitialization by leveraging view-guided warping to ensure view consistency. Through our comprehensive metric framework suitable for novel-view datasets, we show that our method improves view consistency across various diffusion models, demonstrating its broader applicability111project page: https://jwoo-park0.github.io/wave.
@@ 3.2 Adaptive Warp-range Selection
To determine the optimal reference range, we perform the adaptive warp-range selection shown in Fig. 4a. We utilize warped images from 3D warping to compute viewpoint relevance. As the viewpoint shifts, regions that are not visible from the input view become missing regions in the warped images, assigned a value of zero. In order to create warped region masks $M$, we convert warped images into binary regions $W_{1:N}$, where pixels with a value of zero remain 0, and non-zero pixels are set to 1. Then, to measure the relevance between viewpoints (e.g., $i$ and $j$), we compute the IoU $u_{ij}$ on warped region masks to construct the matrix $U=\{u_{ij}\}\in\mathbb{R}^{N\times N}$, which provides a precise measure of their spatial relevance [25]. Based on this matrix, the adaptive range $[\mu_{i}-\sigma_{i},\ \mu_{i}+\sigma_{i}]$ is selected for each viewpoint, for mean $\mu_{i}$ and standard deviation $\sigma_{i}$. Applying the standard deviation of the IoU matrix enables the dynamic determination of reference points, making our approach applicable across various sets of viewpoints.
@@ 3.4 Pose-Aware Noise Initialization
2:$\mathcal{W}^{\prime}\leftarrow\mathcal{W}+\eta\cdot\mathbf{1}_{W=0}$triangleright Add noise $\eta$ to zero regions of $W$
3:$z_{0}\leftarrow\mathcal{E}(W)$ , Sample $\mathbf{\epsilon}\sim\mathcal{N}(0,I)$
4:$z_{T}\leftarrow\text{DDPM Forward}(z_{0},T)$triangleright noise scheduler
5:$F_{T}^{\text{low}}\leftarrow\text{FFT}(z_{T})\odot\mathcal{G(D}_{0})$
6:$F_{T}^{\text{low}}\leftarrow F_{T}^{\text{low}}/\text{mean}(|F_{T}^{\text{low}}|)$triangleright mean of $F_{T}^{\text{low}}$’s magnitude
7:$F_{\epsilon}^{\text{high}}\leftarrow\text{FFT}(\epsilon)\odot(1-\mathcal{G(D}_{0}))$
8:$\epsilon^{\prime}\leftarrow\text{IFFT}(F_{T}^{\text{low}}+F_{\epsilon}^{\text{high}})$
@@ 4.1 Comparison with Baselines
Table 1 compares diffusion-based models with our method. Our method steadily outperforms diffusion-based models across diverse datasets, showing adaptability to diffusion models. Specifically, on the MegaScenes dataset, our method achieves 0.397 in LPIPS-next and 0.920 in CLIPSIM-next, while generating images with higher accuracy in camera viewpoints. VistaDream [42] incorporates a rendering method giving it an advantage in consistency metrics over diffusion models. Thus, VistaDream is compared based on consistency metrics and inference time, unlike other models. Notably, as shown in Fig. 6, the result of MegaScenes+WAVE exhibits significantly lower inference time than VistaDream while achieving comparable performance. This shows that solving the inherent factors of diffusion models can achieve a performance similar to rendering-based methods without excessive computational cost, demonstrating the practicality of our method.
Moreover, to evaluate image quality not captured by our metric framework, we employ the image sequence (i.e., video) dataset, RE10K, to assess the reconstruction and generation capabilities (Table 2). This dataset contains both ground-truth images and camera poses for varying views, allowing the evaluation of image quality. The results show that our method outperforms existing models in both image generation (FID, KID) and reconstruction (PSNR, LPIPS) metrics. The results demonstrate that our method produces consistent images while achieving higher image quality.
@@ 4.3 Ablation Study
Effect of pose-aware noise initialization. Applying pose-aware noise initialization also improves performance over the diffusion model in Table 3. This shows that pose-aware noise initialization effectively reflects view transformations, helping to reduce noise randomness. As shown in Fig. 8, while some objects fail to appear consistently across generated views, color variations do not occur. These results imply that the randomness in noise contributes to color inconsistencies.
@@ 4.4 Downstream Task
We conduct an evaluation by directly using the generated outputs from the diffusion model as inputs to a 3D model to assess rendering performance. In this experiment, VistaDream is excluded since its pipeline already includes 3D rendering. 3D rendering experiments are conducted using 3D Gaussian Splatting [11] with 19 image sets generated by our method and other baseline models [28, 39]. As shown in Table 4, our approach outperforms the original methods. These results indicate that our generated consistent images enhance geometric accuracy and visual coherence in 3D reconstruction and rendering tasks. The potential for future application of our method in diffusion-based 3D rendering tasks is also demonstrated. Results for datasets not included in Table 4 are in the supplementary material due to page limits.
@@ Appendix A Code and Website
The code for WAVE will be released soon. A project website has been created to introduce WAVE in a simple and accessible manner, as well as to showcase a variety of qualitative results and videos. It can be accessed via the following link: project page: https://jwoo-park0.github.io/wave.github.io/
@@ Appendix B More Details of Methods
B.1 Validating the assumption of warp-guided adaptive Attention
We propose warp-guided adaptive attention to manipulate attention, utilizing warped region masks. This approach is based on the hypothesis that the attention mechanism in the U-Net decoder layers preserves spatial position correspondence. Leveraging this assumption, we design warped region masks that align with the attention maps, enabling direct attention manipulation. This method efficiently incorporates the novel viewpoint information into the attention mechanism, contributing to generating consistent images.
@@ C.1 Video consistency metrics
[…]
@@ D.1 Implementation details
For the novel view synthesis diffusion model, this work utilizes pre-trained models such as ZeroNVS [28] and MegaScenes [39], which generate images at a resolution of 256×256, maintained across all experiments. DDIM with 50 inference steps is adopted for sampling. Additionally, the attention map dropout technique from Tewel et al. [38] is implemented, using a fixed dropout ratio of 0.2, with no significant performance changes observed when varying the dropout ratio. For noise initialization, a Gaussian low-pass filter is applied, following the filter used in prior studies [44, 24]. During the noise application to latent variables, the noise level is fixed at 950 after the latent variables pass through the VAE encoder. The warp algorithm is implemented using the Pyrender library, equal to MegaScenes. Camera parameters are also configured in the OpenGL environment, as required by Pyrender. For warping, the depth map is extracted using the Depth Anything model [45], a monocular depth estimation model, following the approach of the previous work, MegaScenes.
@@ D.5 Downstream task
D.6 Quantitative evaluation of adaptive warp-range selection effect
Table 7: Camera accuracy in warp-guided adaptive attention. We show quantitative results to evaluate whether warp-guided adaptive attention improves viewpoint accuracy. Our method, warp-guided adaptive attention, achieves higher camera pose accuracy compared to batch self-attention, which has been used in previous studies.
@@ D.6 Quantitative evaluation of adaptive warp-range selection effect
MegaScenes + conventional batch attention 0.412 0.299 17.13
@@
MegaScenes + conventional batch attention 0.383 0.281 16.09
@@
MegaScenes + conventional batch attention 0.321 0.231 13.25
@@
Appendix E More Qualitative Results
E.1 Camera parameter visualization
@@ E.2 Additional samples
Appendix F Failure Cases

### v2

@@ Abstract
Generating high-quality novel views of a scene from a single image requires maintaining structural coherence across different views, referred to as view consistency. While diffusion models have driven advancements in novel view synthesis, they still struggle to preserve spatial continuity across views. Diffusion models have been combined with 3D models to address the issue, but such approaches lack efficiency due to their complex multi-step pipelines. This paper proposes a novel view-consistent image generation method which utilizes diffusion models without additional modules. Our key idea is to enhance diffusion models with a training-free method that enables adaptive attention manipulation and noise reinitialization by leveraging view-guided warping to ensure view consistency. Through our comprehensive metric framework suitable for novel-view datasets, we show that our method improves view consistency across various diffusion models, demonstrating its broader applicability  [^1] Project page: jwoo-park0.github.io/wave
@@ 3.2 Adaptive Warp-range Selection
To determine the optimal reference range, we perform the adaptive warp-range selection shown in Fig. 4a. We utilize warped images from 3D warping to compute viewpoint relevance. As the viewpoint shifts, regions that are not visible from the input view become missing regions in the warped images, assigned a value of zero. In order to create warped region masks $M$, we convert warped images into binary regions $W_{1:N}$, where pixels with a value of zero remain 0, and non-zero pixels are set to 1. Then, to measure the relevance between viewpoints (e.g., $i$ and $j$), we compute the IoU $u_{ij}$ on warped region masks to construct the matrix $U=\{u_{ij}\}\in\mathbb{R}^{N\times N}$, which provides a precise measure of their spatial relevance [25]. Based on this matrix, the adaptive range $\mu_{i}-\sigma_{i},\ \mu_{i}+\sigma_{i}]$ is selected for each viewpoint, for mean $\mu_{i}$ and standard deviation $\sigma_{i}$. Applying the standard deviation of the IoU matrix enables the dynamic determination of reference points, making our approach applicable across various sets of viewpoints.
@@ 3.4 Pose-Aware Noise Initialization
[…]
@@ 4.1 Comparison with Baselines
Table 1 compares novel view synthesis diffusion models with our method. Our method steadily outperforms diffusion-based models across diverse datasets, showing adaptability to diffusion models. Specifically, on the MegaScenes dataset, our method achieves 0.397 in LPIPS-next and 0.920 in CLIPSIM-next, while generating images with higher accuracy in camera viewpoints. VistaDream[42] and Viewcrafter[47] incorporate a rendering method giving them an advantage in consistency metrics over diffusion models. Thus, these models are compared based on consistency metrics and inference time, unlike other models. Additionally, we include Genwarp [12] as a baseline to provide a more comprehensive comparison. Notably, as shown in Fig. 6, the result of MegaScenes+WAVE exhibits significantly lower inference time than rendering-based methods while achieving comparable performance. This shows that solving the inherent factors of diffusion models can achieve a performance similar to rendering-based methods without excessive computational cost, demonstrating the practicality of our method.
Moreover, to evaluate image quality not captured by our metric framework, we employ the image sequence (i.e., video) dataset, RE10K, to assess the reconstruction and generation capabilities, as illustrated in Table 2. This dataset contains both ground-truth images and camera poses for varying views, allowing the evaluation of image quality. The results show that our method outperforms existing models in both image generation (FID, KID) and reconstruction (PSNR, LPIPS) metrics. The results demonstrate that our method produces consistent images while achieving higher image quality.
@@ 4.3 Ablation Study
Effect of pose-aware noise initialization. Applying pose-aware noise initialization also improves performance over the diffusion model in Table 8. This shows that pose-aware noise initialization effectively reflects view transformations, helping to reduce noise randomness. As shown in Fig. 8, while objects fail to appear across generated views, color variations do not occur. These results imply that the randomness in noise contributes to color inconsistencies.
@@ 4.4 Downstream Task
We conduct an evaluation by directly using the generated outputs from the diffusion model as inputs to a 3D model to assess rendering performance. In this experiment, VistaDream is excluded since its pipeline already includes 3D rendering. 3D rendering experiments are conducted using 3D Gaussian Splatting [11] with 19 image sets generated by our method and other baseline models [28, 39]. As shown in Table 4, our approach outperforms the original methods. These results indicate that our generated consistent images enhance geometric accuracy and visual coherence in 3D reconstruction and rendering tasks. The potential for future application of our method in diffusion-based 3D rendering tasks is also demonstrated. Results for datasets not included in Table 4 are in the supplementary material.
@@
Acknowledgement. This work was supported in part by IITP RS-2024-00457882 (National AI Research Hub Project), IITP 2020-0-01361, NRF RS-2024-00345806, NRF RS-2023-00219019, G21004687962, and RS-2024-00403860 (Korea Basic Science Institute, National Research Facilities and Equipment Center).
@@ Appendix A Code and Website
A project website has been created to introduce WAVE in a simple and accessible manner, as well as to showcase a variety of qualitative results, implementation codes and videos. It can be accessed via the following link: project page: https://jwoo-park0.github.io/wave.github.io/
@@ Appendix B More Details of Methods
B.1 Validating the assumption of WGAA
We propose warp-guided adaptive attention, WGAA, to manipulate attention, utilizing warped region masks. This approach is based on the hypothesis that the attention mechanism in the U-Net decoder layers preserves spatial position correspondence. Leveraging this assumption, we design warped region masks that align with the attention maps, enabling direct attention manipulation. This method efficiently incorporates the novel viewpoint information into the attention mechanism, contributing to generating consistent images.
@@ C.1 Video consistency metrics
where: $\psi(I)$ represents the CLIP embedding of image $I$, $\cdot$ denotes the dot product, $\|\cdot\|$ represents the Euclidean norm. As illustrated in Fig. 10, LPIPS-first and CLIPSIM-first calculate LPIPS and CLIP similarity between the input viewpoint and all other viewpoints. In contrast, LPIPS-next and CLIPSIM-next measure consistency between adjacent viewpoints. Since the Next metrics are less affected by viewpoint changes compared to the First metrics, they provide a more effective measure of view consistency.
@@ D.1 Implementation details
For the novel view synthesis diffusion model, this work utilizes pre-trained models such as ZeroNVS [28] and MegaScenes [39], which generate images at a resolution of 256×256, maintained across all experiments. DDIM with 50 inference steps is adopted for sampling. Additionally, the attention map dropout technique from Tewel et al. [38] is implemented, using a fixed dropout ratio of 0.2, with no significant performance changes observed when varying the dropout ratio. For noise initialization, a Gaussian low-pass filter is applied, following the filter used in prior studies [44, 24]. During the noise application to latent variables, the noise level is fixed at 950 after the latent variables pass through the VAE encoder. The warp algorithm is implemented using the Pyrender library, equal to MegaScenes. Camera parameters are also configured in the OpenGL environment, as required by Pyrender. For warping, the depth map is extracted using the Depth Anything model [45], a monocular depth estimation model, following the approach of the previous work, MegaScenes. All experiments are conducted using an RTX A6000 GPU.
@@ D.5 Downstream task
Table 7: Camera accuracy in warp-guided adaptive attention. We show quantitative results to evaluate whether warp-guided adaptive attention improves viewpoint accuracy. Our method, warp-guided adaptive attention, achieves higher camera pose accuracy compared to batch self-attention (conventional_att), which has been used in previous studies.
@@
MegaScenes + conventional_att 0.412 0.299 17.13
@@
MegaScenes + conventional_att 0.383 0.281 16.09
@@
MegaScenes + conventional_att 0.321 0.231 13.25
@@
Appendix E Additional Experiments
E.1 Effect of WGAA on Camera Accuracy
E.2 Analysis by Depth Map Quality
We conduct additional experiments to investigate its impact on view consistency. As part of this, we evaluate the performance across different Depth Anything [45] model sizes, where smaller models show degraded performance (AbsRel). The table indicates that although the depth metric degrades, view consistency shows little variation across different depth estimation models, outperforming base MegaScenes. This demonstrates the robustness of our method.
Table 8: Performance by Depth Map Quality. We evaluate the performance of our method when applied to different model sizes of Depth Anything (DA), and report the corresponding quantitative results.
Method AbsRel $\downarrow$ Next-LPIPS $\downarrow$ Next-CLIPSIM $\uparrow$ First-LPIPS $\downarrow$ First-CLIPSIM $\uparrow$
MegaScenes - 0.607 0.861 0.609 0.813
WAVE (DA-S) 0.053 0.415 (-0.192) 0.926 (+0.065) 0.568 (-0.041) 0.796 (-0.017)
WAVE (DA-B) 0.046 0.399 (-0.208) 0.927 (+0.066) 0.558 (-0.051) 0.804 (-0.009)
WAVE (DA-L) 0.043 0.397 (-0.210) 0.920 (+0.059) 0.561 (-0.048) 0.826 (+0.013)
E.3 More Comparisons with PANI
We conduct an additional experiment to evaluate PANI by comparing it with other training-free noise initialization methods designed for consistency. Specifically, we compare our method with FrameInit (3D FFT) from ConsistI2V [24]. Unfortunately, MultiDiff [21] is excluded from the comparison due to the unavailability of its code. This experiment is conducted on the MegaScenes dataset. The results show that PANI outperforms in both view consistency and camera pose accuracy. Unlike FrameInit method for consistent video, which does not incorporate view information, our method integrates such information, enabling the generation of view-consistent images.
Table 9: Comparison with other noise initialization method. Additional experiments are conducted to validate the effectiveness of PANI, comparing it with other noise initialization methods proposed in previous studies. Evaluation is based on both consistency and camera pose accuracy.
Method Next-LPIPS $\downarrow$ Next-CLIPSIM $\uparrow$ First-LPIPS $\downarrow$ First-CLIPSIM $\uparrow$
WAVE w/ FrameInit 0.433 0.919 0.610 0.813
WAVE w/ PANI 0.397 0.920 0.561 0.826
Method Frobenius Norm (Rotation) Rotation Angle Difference Angular Consistency
WAVE w/ FrameInit 0.402 0.291 16.67
WAVE w/ PANI 0.382 0.277 15.91
Appendix F More Qualitative Results
F.1 Camera parameter visualization
@@ F.2 Additional samples
Appendix G Failure Cases